The 2.7 kg cylinder B and 1.8 kg wedge A are released from a bigger triangular wedge A wich is fixed. the angle of inclination of A is 30 degrees frm the horizontal.Assuming that cylinder rolls without sliding on the wedge and neglecting the fricyion between a and c determine
(a) the acceleration of wedge
(b) the angular acceleration of cylinder
(g=10m/s^2)




Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Aditya, 

acceleration of wedge=Mwgsin(30)/M=5m/s^2

a=linear acceleration,A=angular acceleration,f=friction,r=radius of cylinder,m=mass

a=rA, ma=mgsin(30)-f,fr=(1/2)mr2A, solving the three simultaneous equations we get a=(2/3)gsin(30)=10/3 m/s^2

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras